# QEA: An Accelerator for Quantum Circuit Simulation with Resources Efficiency and Flexibility

### v2

@@ I Introduction
To address these problems, we offer a hardware architecture named Quantum Emulation Accelerator (QEA) intended to support a wide range of applications with a high #Qubits while maintaining both flexibility and high performance at low costs. Our approach combines hardware and software enhancements to efficiently simulate quantum circuits. On the software side, we still follow the state-vector simulator but enhance the performance of the overall process via gate-group-and-fusion and compact tensor product (TP) techniques. On the hardware side, we propose four key approaches, including:
@@ IV-B Benchmarking setting
To prove the generality of QEA, we evaluated 19 circuit templates, each labeled with a circuit ID from 1 to 19. The full circuit representation is presented in Fig.  3 at [15]. All circuits are constructed from a combination of topologies, including chain, alternating, all-to-all [16], and rotation; as shown in Fig. 3. In circuit design, the chain topology arranges qubits in a linear sequence, where each qubit is coupled only to its immediate neighbors. The alternating topology extends this by introducing a bipartite structure, connecting qubits in two subsets with edges between them. In contrast, the all-to-all topology allows every qubit to interact directly with every other, forming a complete graph. The final, rotation $i$ takes a role as parameter holder and provides no entanglement.
@@ 2 $\times$ Intel Stratix 10
This section compared QEA to other related studies on FPGA to provide a more comprehensive evaluation of performance and efficiency. The QFT technique was utilized to compare with other studies since it is widely used in the creation of hardware designs. Table I compares QEA’s hardware results to other FPGA-based works [9, 5, 6, 7, 4] in terms of frequency, reconfigurable, #Qubits, and normalized gate speed due to differences in supported qubits, precision, hardware resources, and the number of gates in quantum circuits.
@@ V Conclusion
In summary, this paper introduced QEA, a quantum emulator, to address current challenges in achieving high flexibility, memory efficiency, low cost, and high performance. QEA was presented with four key ideas: optimized memory allocation management, open PE, flexible ALU, and simplified CX swapper. The complete verification and evaluation confirmed the exceptional properties of QEA in terms of low MSE, high fidelity, less memory usage, short execution time, and high flexibility compared to other comparable works in supporting a variety of applications with varying #Qubits from 3 to 17. In future work, QEA can be enhanced and integrated into a system using external data memories and multiple QEA cores to get higher performance with higher supported #Qubits.

### v3

@@ I Introduction
To address these problems, we offer a hardware architecture named Quantum Emulation Accelerator (QEA) intended to support a wide range of applications with a high #Qubits while maintaining both flexibility and high performance at low hardware resources. Our approach combines hardware and software enhancements to efficiently simulate quantum circuits. On the software side, we still follow the state-vector simulator but enhance the performance of the overall process via gate-group-and-fusion and compact tensor product (TP) techniques. On the hardware side, we propose four key approaches, including:
@@ IV-B Benchmarking setting
To prove the generality of QEA, we evaluated 19 circuit templates [15], each labeled with a circuit ID from 1 to 19. The full circuit representation is presented in Fig.  3. All circuits are constructed from a combination of topologies, including chain, alternating, all-to-all [16], and rotation; as shown in Fig. 3. In circuit design, the chain topology arranges qubits in a linear sequence, where each qubit is coupled only to its immediate neighbors. The alternating topology extends this by introducing a bipartite structure, connecting qubits in two subsets with edges between them. In contrast, the all-to-all topology allows every qubit to interact directly with every other, forming a complete graph. The final, rotation $i$ takes a role as parameter holder and provides no entanglement.
@@ 2 $\times$ Intel Stratix 10
This section compares QEA to other related studies on FPGA to provide a more comprehensive evaluation of performance and efficiency. The QFT technique was utilized to compare with other studies since it is widely used in the creation of hardware designs. Table I compares QEA’s hardware results to other FPGA-based works [9, 5, 6, 7, 4] in terms of frequency, reconfigurable, #Qubits, and normalized gate speed due to differences in supported qubits, precision, hardware resources, and the number of gates in quantum circuits.
@@ V Conclusion
In summary, this paper introduced QEA, a quantum emulator, to address current challenges in achieving high flexibility, memory efficiency, low hardware resources, and high performance. QEA was presented with four key ideas: optimized memory allocation management, open PE, flexible ALU, and simplified CX swapper. The complete verification and evaluation confirmed the exceptional properties of QEA in terms of low MSE, high fidelity, less memory usage, short execution time, and high flexibility compared to other comparable works in supporting a variety of applications with varying #Qubits from 3 to 17. In future work, QEA can be enhanced and integrated into a system using external data memories and multiple QEA cores to get higher performance with higher supported #Qubits.